# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ Appendix C Concentration Bounds
The value in an asymptotic equicontinuity expansions is that we may often be able to assume that the expected function $f_{n}(p)$ is differentiable even when the $F_{in}(p)$ themselves are not (e.g., if the $F_{in}(p)$ capture choices regarding supply and demand that may vary discontinuously in prices). Asymptotic equicontinuity then allows us to approximate sample averages of $F_{in}(p)$ by Taylor expanding $f_{n}(p)$.
@@
Our approach to establishing asymptotic equicontinuity of relevant quantities in our marketplace model starts by bounding the bracketing number of approximately monotone functions. Let $F_{i}:\mathbb{R}^{J}\rightarrow\mathbb{R}$ be random functions belonging to some set $\mathcal{F}$, and let $\mathcal{S}\subset\mathbb{R}^{J}$ be the set of possible market equilibrium prices. For any pair $p_{-},\,p_{+}$ in $\mathbb{R}^{J}$, we define the $\mathcal{F}$-bracket over $\mathcal{S}$ as
$[p_{-},\,p_{+}]_{\mathcal{F}}=\left\{p\in\mathcal{S}:F(p_{-})\geq F(p)\geq F(p_{+})\text{ for all }F\in\mathcal{F}\right\}.$ (74)
Given a distribution $P$ over the random functions, i.e., $F_{i}\,{\buildrel\text{iid}\over{\sim}\,}P$, we define the $L_{2}(P)$-length of the bracket as $d^{2}_{P}(p_{-},\,d_{+})=\mathbb{E}_{P}\left[(F_{i}(p_{-})-F_{i}(p_{+}))^{2}\right]$, and the $\varepsilon$-bracketing number of $\mathcal{S}$ under $P$ as
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))=\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}[p_{-}(k),\,p_{+}(k)]_{\mathcal{F}}\text{ with }d_{P}\left(p_{-}(k),\,p_{+}(k)\right)\leq\varepsilon\text{ for all }k\right\}.$ (75)
@@
Consider a class of random functions $F_{i}:\mathbb{R}^{J}\rightarrow\mathbb{R}$ with $F_{i}\in\mathcal{F}$ almost surely, and let $\mathcal{S}$ is a compact subset of $\mathbb{R}^{J}$. Suppose that, for all $p\in\mathbb{R}^{J}$, $\varepsilon\leq 1$ and $\left\lVert\delta\right\rVert_{2}\leq C\varepsilon$, we have
$F_{i}(p-\varepsilon e_{j})\geq F_{i}(p+\delta)\geq F_{i}(p+\varepsilon e_{j}).$ (76)
Suppose also that the $F_{i}$ have a distribution $P$ under which
$\mathbb{E}_{P}\left[\left(F_{i}(p)-F_{i}(p^{\prime})\right)^{2}\right]\leq L\left\lVert p-p^{\prime}\right\rVert_{2}\text{ for all }p,\,p^{\prime}\in\mathbb{R}^{J}.$ (77)
@@
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))\leq\theta(\mathcal{S},\,C,\,L)\,\varepsilon^{-2J}.$ (78)
@@
$B_{C\zeta}(p)\subseteq[p-\zeta e_{j},\,p+\zeta e_{j}]_{\mathcal{F}}.$
Meanwhile, by (77) we have $d_{P}^{2}\left(p-\zeta e_{j},\,p+\zeta e_{j}\right)\leq 2L\zeta$, and so $d_{P}\left(p-\zeta e_{j},\,p+\zeta e_{j}\right)\leq\varepsilon$ if $\zeta=\varepsilon^{2}/(2L)$. These two facts together imply that
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))\leq\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{C(\varepsilon^{2}/(2L))}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\},$
@@
We say that a triangular array of random functions $F_{in}$ is $\zeta$-subgaussian over $\mathcal{S}$ if there exists a constant $C$ such that, for all $n$ and $t\geq 1$,
@@
Under Assumptions 3, 4 and 5, let $U_{i}$ be random price perturbations with $\mathbb{E}\left[U_{i}\right]=0$ and $\left\lvert U_{i}\right\rvert\leq h_{n}$ almost surely for some sequence $h_{n}\rightarrow 0$. Suppose furthermore that $W_{i}$ is Bernoulli-randomized as in Design 1. Then, the following function classes are $\zeta$-subgaussian with $\zeta=2J$:
@@
The result for $Z_{i}(w,\,p+U_{i})$ and $Z_{i}(W_{i},\,p+U_{i})$ follows immediately from the bracketing bounds established in Lemma 18, boundedness as required in Assumption 3, and Theorem 2.14.9 of van der Vaart & Wellner (1996). The result for $W_{i}/\pi_{i}\,Z_{i}(W_{i},\,p+U_{i})$ and $(1-W_{i})/(1-\pi_{i})\,Z_{i}(W_{i},\,p+U_{i})$, and thus also that of their difference, follows by the same argument because $W_{i}/\pi_{i}$ and $(1-W_{i})/(1-\pi_{i})$ are positive and bounded by $\eta^{-1}$, and thus these functions satisfy the conditions of Lemma 18 with a constant $L/\eta$ in (77).
Next recall to functions constructed in therms of $Y_{i}(w,\,p)$, recall that by Assumption 3 there is a random function $H_{i}(w,\,p)$ that is Lipschitz continuous in $p$ such that
$Y_{i}(w,\,p+U_{i})$ $=H_{i}(w,\,p+U_{i})+\left(\alpha+\beta^{\top}(p+U_{i})\right)\cdot Z_{i}(w,\ p+U_{i})$ (81)
$=H_{i}(w,\,p+U_{i})+\alpha^{\top}Z_{i}(w,\ p+U_{i})+p^{\top}\beta\,Z_{i}(w,\ p+U_{i})+U_{i}^{\top}\beta Z_{i}(w,\ p+U_{i}).$
Our above concentration result immediately generalizes to the 2nd and 3rd terms above, because the multipliers are non-random and $p$ takes values within a compact set $\mathcal{S}$. To control the 4th term, we note that
$U_{i}^{\top}\beta Z_{i}(w,\ p+U_{i})=\sum_{j,\,k=1}^{J}\left(U_{ij}\beta_{jk}\right)_{+}Z_{ik}(w,\ p+U_{i})-\sum_{j,\,k=1}^{J}\left(-U_{ij}\beta_{jk}\right)_{+}Z_{ik}(w,\ p+U_{i}),$
i.e., the term can be written as $2J^{2}$ weighted averages of the net demand for a given good, with non-negative, bounded weights. We can thus obtain concentration for this term by applying our above argument a finite number of times.
Finally, to control the contribution of the $H_{i}(w,\,p)$ term, we note that bracketing bounds for Lipschitz function classes are standard; see, e.g., Section 2.7.4 of van der Vaart & Wellner (1996). Adapting these results to our setting, let
@@
Clearly, for any $p\in\mathbb{R}^{J}$,
@@
and $d_{P}((p,\,-M\varepsilon),\,(p,\,M\varepsilon))=2M\varepsilon$ for any distribution $P$ over the functions $A(\cdot)$. Thus, using notation from Lemma 18,
$N_{[]}(\varepsilon,\,\mathcal{S},\,L_{2}(P))\leq\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{\varepsilon/(2M)}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\},$ (82)
which is in turn $O(\varepsilon^{-J})$ by the same argument as used in (79). This bracketing bound can be applied to $H_{i}(w,\,p+U_{i})$, $H_{i}(W_{i},\,p+U_{i})$, etc., thus enabling us to again apply Theorem 2.14.9 of van der Vaart & Wellner (1996). Other functions constructed using $Y_{i}(w,\,p)$ can be controlled analogously. ∎
Lemma 20.
Under the setting of Lemma 19, the following function classes are asymptotically equicontinuous, i.e., they satisfy (73) at all $p\in\mathcal{S}$:
@@
Furthermore, the residuals in the asymptotic equicontinuity expansion,
satisfy $\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]=0$.
@@
We start with the functions constructed using $Z_{i}(w,\,p)$. Fix a good $j\in\left\{1,\,\ldots,\,J\right\}$. By Assumption 3, we see that $Z_{ij}(w,\,p+U_{i})$ and $Z_{ij}(W_{i},\,p+U_{i})$ satisfy the conditions of Lemma 18; and, by weak continuity of $Z_{ij}(w,\,p)$, we have
$\lim_{n\rightarrow\infty}\operatorname{Cov}\left[Z_{ij}(w,\,p+U_{i}),\ Z_{ij}(w,\,p^{\prime}+U_{i})\right]=\operatorname{Cov}\left[Z_{ij}(w,\,p),\ Z_{ij}(w,\,p^{\prime})\right].$ (84)
Thus, for both $F_{in}(p)=Z_{ij}(w,\,p+U_{i})$ and $F_{in}(p)=Z_{ij}(W_{i},\,p+U_{i})$, weak convergence of the empirical process
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(F_{in}(p)-f_{n}(p)\right)$
to a Gaussian process indexed by $p$ follows from Lemma 18 and Theorem 2.11.23 of van der Vaart & Wellner (1996). Furthermore, asymptotic equicontinuity follows by weak continuity as required in Assumption 5, as weak continuity implies that the limiting Gaussian process has continuous sample paths.
Asymptotic equicontinuity of $(W_{i}/\pi_{i}-(1-W_{i})/(1-\pi_{i}))Z_{ij}(W_{i},\,p+U_{i})$ follows analogously via the argument made in Lemma 19. We have thus shown asymptotic equicontinuity separately for the $j$-th entry of each function class constructed using $Z_{i}(w,\,p)$; and this implies asymptotic equicontinuity of the full (vector-valued) functions themselves also.
@@
Now, pick any $c>0$. By asymptotic equicontinuity, $\limsup_{n\rightarrow\infty}\mathbb{P}\left[\left\lvert R_{n}\right\rvert>c/\sqrt{n}\right]=0$. Furthermore
@@
where for the second inequality we used Cauchy-Schwarz. Since this holds for any $c>0$, we see that in fact $\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]=0$.
To establish asymptotic equicontinuity of the functions constructed using $Y_{i}(w,\ p)$, we again leverage the decomposition (81), and verify that each term is asymptotically equicontinuous on its own. We again focus on deriving asymptotic equicontinuity for $F_{i}(p)=Y_{i}(w,\ p+U_{i})$; the extension to other functions is immediate. For the first term in (81), our Lipschitz assumption yields
$\lim_{n\rightarrow\infty}\operatorname{Cov}\left[H_{i}(w,\,p+U_{i}),\ H_{i}(w,\,p^{\prime}+U_{i})\right]=\operatorname{Cov}\left[H_{i}(w,\,p),\ H_{i}(w,\,p^{\prime})\right],$ (85)
and so we can use (82) and Theorem 2.11.23 of van der Vaart & Wellner (1996) to establish weak convergence to a Gaussian process; and continuity of the limiting process is immediate from Lipschitz continuity of $H_{i}(w,\,p)$. Meanwhile, asymptotic equicontinuity of the 2nd, 3rd and 4th terms follows by adapting arguments already used to study net demand, and discussed in the proof of Lemma 19. ∎

### v4

@@ Appendix C Concentration Bounds
The motivation behind establishing asymptotic equicontinuity expansions is that we may often be able to assume that the expected function $f_{n}(p)$ is differentiable even when the $F_{in}(p)$ themselves are not (e.g., if the $F_{in}(p)$ capture choices regarding supply and demand that may vary discontinuously in prices). Asymptotic equicontinuity then allows us to approximate sample averages of $F_{in}(p)$ by Taylor expanding $f_{n}(p)$.
@@
Our approach to establishing asymptotic equicontinuity of relevant quantities in our marketplace model starts by bounding the bracketing number of approximately monotone functions. Let $F_{in}:\mathbb{R}^{J}\rightarrow\mathbb{R}$ be random functions drawn i.i.d. from a distribution $Q_{n}$, and let $\mathcal{S}\subset\mathbb{R}^{J}$ be the set of possible market equilibrium prices. We are interested in a bracket for the function class $\mathcal{F}_{n}=\{F_{in}(\cdot)\to F_{in}(p):p\in\mathcal{S}\}$. For any pair $p_{-},\,p_{+}$ in $\mathbb{R}^{J}$, we define a bracket as
$[p_{-},\,p_{+}]=\left\{p\in\mathcal{S}:F_{in}(p_{-})\geq F_{in}(p)\geq F_{in}(p_{+})\text{ for all }F_{in}(\cdot)\right\}.$ (81)
We define the square of the $L_{2}(Q_{n})$-length of the bracket as $d^{2}_{Q_{n}}(p_{-},\,p_{+})=\mathbb{E}_{Q_{n}}\left[(F_{in}(p_{-})-F_{in}(p_{+}))^{2}\right]$, and the $\varepsilon$-bracketing number of $\mathcal{F}_{n}$ under $Q_{n}$ as
$N_{[]}(\varepsilon,\,\mathcal{F}_{n},\,L_{2}(Q_{n}))=\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}[p_{-}(k),\,p_{+}(k)]\text{ with }d_{Q_{n}}\left(p_{-}(k),\,p_{+}(k)\right)\leq\varepsilon\text{ for all }k\right\}.$ (82)
@@
Consider a class of random functions $F_{in}:\mathbb{R}^{J}\rightarrow\mathbb{R}$ with $F_{in}\in\mathcal{G}_{n}$ almost surely, and $\mathcal{S}$ is a compact subset of $\mathbb{R}^{J}$. Let $\mathcal{F}_{n}=\{F_{in}(\cdot)\to F_{in}(p):p\in\mathcal{S}\}$. Suppose that, for all $p\in\mathcal{S}$, $\varepsilon\leq 1$ and $\left\lVert\delta\right\rVert_{2}\leq C\varepsilon$, we have
$F_{in}(p-\varepsilon e_{j})\geq F_{in}(p+\delta)\geq F_{in}(p+\varepsilon e_{j}).$ (83)
Suppose also that the $F_{in}$ have a distribution $Q_{n}$ under which
$\mathbb{E}_{Q_{n}}\left[\left(F_{in}(p)-F_{in}(p^{\prime})\right)^{2}\right]\leq L\left\lVert p-p^{\prime}\right\rVert_{2}\text{ for all }p,\,p^{\prime}\in\mathbb{R}^{J}.$ (84)
@@
$N_{[]}(\varepsilon,\,\mathcal{F}_{n},\,L_{2}(Q_{n}))\leq\theta(\mathcal{S},\,C,\,L)\,\varepsilon^{-2J}.$ (85)
@@
$B_{C\zeta}(p)\subseteq[p-\zeta e_{j},\,p+\zeta e_{j}].$
Meanwhile, by (84) we have $d_{Q_{n}}^{2}\left(p-\zeta e_{j},\,p+\zeta e_{j}\right)\leq 2L\zeta$, and so $d_{Q_{n}}\left(p-\zeta e_{j},\,p+\zeta e_{j}\right)\leq\varepsilon$ if $\zeta=\varepsilon^{2}/(2L)$. These two facts together imply that
$N_{[]}(\varepsilon,\,\mathcal{F}_{n},\,L_{2}(Q_{n}))\leq\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{C(\varepsilon^{2}/(2L))}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\},$
@@
We say that a random function $F_{in}$ is $\zeta$-subgaussian over $\mathcal{S}$ if there exists a constant $C$ such that, for all $n\geq 1$ and $t>0$, that
@@
Under Assumptions 4, 5 and 6, let $U_{i}$ be random price perturbations with $\mathbb{E}\left[U_{i}\right]=0$ and $\left\lvert U_{i}\right\rvert\leq h_{n}$ almost surely for some sequence $h_{n}\rightarrow 0$. Suppose furthermore that $W_{i}$ is Bernoulli-randomized as in Design 1. Then, the following random functions are $\zeta$-subgaussian:
@@
For $Z_{i}(w,\,p+U_{i})$ and $Z_{i}(W_{i},\,p+U_{i})$, we can use the bracketing bounds in Lemma 20, which imply $N_{[]}(\varepsilon,\mathcal{F}_{n},L_{2}(Q_{n}))\leq({K}/{\varepsilon})^{\zeta}$, for $\zeta=2J$, and $\mathcal{F}_{n}$ is constructed by varying $Z_{i}(w,\,p+U_{i})$ or $Z_{i}(W_{i},\,p+U_{i})$ over $p\in\mathcal{S}$. The tail bound follows from Theorem 2.14.9 of van der Vaart & Wellner (1996), given boundedness of the net demand functions and the polynomial bound on the $\varepsilon$-bracketing number. The result for $W_{i}/\pi_{i}\,Z_{i}(W_{i},\,p+U_{i})$, $(1-W_{i})/(1-\pi_{i})\,Z_{i}(W_{i},\,p+U_{i})$ and their differences follows by the same argument because $W_{i}/\pi_{i}$ and $(1-W_{i})/(1-\pi_{i})$ are positive and bounded by $\eta^{-1}$, and thus these functions satisfy the conditions of Lemma 20 with a constant $L/\eta$ in (84).
For outcomes, we have the decomposition $Y_{i}(w,p)=H_{i}(w,p)+\psi(\Gamma_{i}(w),Z_{i}(w,p),p)$. We first work with $H_{i}(w,p)$. Note that bracketing bounds for Lipschitz function classes are standard; see, e.g., Section 2.7.4 of van der Vaart & Wellner (1996). Adapting these results to our setting, we let $F_{i}(w,p)=H_{i}(w,p)+c$ and =$\mathcal{F}=\{F_{i}(w,\cdot)\to F_{i}(w,p):p\in\mathcal{S}\}$.
By the Lipschitz property, for any $p\in\mathbb{R}^{J}$,
@@
and $d_{P}((p,\,-M\varepsilon),\,(p,\,M\varepsilon))\leq 2M\varepsilon$ for any distribution $Q$ over the functions $H_{i}(w,p)$.
Thus, using notation from Lemma 20,
$N_{[]}(\varepsilon,\,\mathcal{F},\,L_{2}(Q))\leq\inf\left\{K:\mathcal{S}\subseteq\bigcup_{k=1}^{K}B_{\varepsilon/(2M)}(p_{k})\text{ with }p_{k}\in\mathbb{R}^{J}\right\},$ (88)
which is in turn $O(\varepsilon^{-J})$ by the same argument as used in (86). This bracketing bound can be applied to $H_{i}(w,\,p+U_{i})$, $H_{i}(W_{i},\,p+U_{i})$, etc., thus enabling us to again apply Theorem 2.14.9 of van der Vaart & Wellner (1996).
For the second component of the outcome function, it is useful to use a covering number bound rather than a bracketing number bound. For a function class $\mathcal{F}_{n}$, we can define the $\varepsilon$-covering number of $\mathcal{F}_{n}$ under a distribution $Q_{n}$ as $N(\varepsilon,\mathcal{F}_{n},L_{2}(Q_{n}))$. This is defined as the minimum number of balls of radius $\varepsilon$ in $L_{2}(Q_{n})$ norm required to cover $\mathcal{F}_{n}$. $\varepsilon$-covering numbers are bounded by $2\varepsilon$-bracketing numbers, as shown in Section 2.1.1 of van der Vaart & Wellner (1996). Theorem 2.14.9 of van der Vaart & Wellner (1996) provides a tail bound as in the Lemma statement for function classes with
$\sup\limits_{Q_{n}}N(\varepsilon,\mathcal{F}_{n},L_{2}(Q_{n}))\leq\left(\frac{K}{\varepsilon}\right)^{\zeta}$
for constants $K,\zeta>0$. It remains to provide such a bound for the class $\{Y_{i}(\cdot)\to Y_{i}(w,p):p\in\mathcal{S}\}$.
Let $G_{in}=\phi(F_{i1n},\ldots,F_{iMn})$ be a composition of $M$ random functions, drawn from distribution $Q_{n}$. Let $\mathcal{G}_{n}=\{\phi(F_{i1n}(\cdot),\ldots,F_{iMn}(\cdot))\to\phi(F_{i1n}(p),\ldots,F_{iMn}(p)):p\in\mathcal{S}\}.$ and $\mathcal{F}_{mn}=\{F_{imn}(\cdot)\to F_{imn}(p):p\in\mathcal{S}\}$. Lemma A.6 of Chernozhukov et al. (2014) indicates that as long as $\phi(\cdot)$ is Lipschitz in each of its $M$ arguments, then if there is a polynomial bound on the $\varepsilon$-covering number of $\mathcal{F}_{mn}$ for $m\in\{1,\ldots,M\}$, there is also a polynomial bound on the $\varepsilon$-covering number of $\mathcal{G}$. Recall that a $\varepsilon$-covering number is bounded by a $2\varepsilon$-bracketing number and our bracketing numbers do not depend on $Q_{n}$. We can now apply the composition result from Chernozhukov et al. (2014) to finish the proof for $Y_{i}(w,p)$, since we already showed that $H_{i}(w,p)$ has the right kind of bound on the bracketing number of its function class, $\psi(\cdot)$ is Lipschitz in each of its arguments, and each of its arguments are functions of $p$ that have a polynomial bound on their $\varepsilon$-covering numbers. ∎
Lemma 22.
@@
$\lim_{n\rightarrow\infty}\operatorname{Cov}\left[F_{in}(p),\ F_{in}(p^{\prime})\right]=\Sigma(p,\,p^{\prime})\text{ for all }p,\,p^{\prime}.$ (89)
Then, averages of that function are asymptotically equicontinuous, i.e., they satisfy (80) at all $p\in\mathcal{S}$. Furthermore, the residuals in the asymptotic equicontinuty expansion,
satisfy $\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]=0$. In particular, under the assumptions of Lemma 21, averages of the following random functions are asymptotically equicontinuous and have residuals in the asymptotic equicontinuity expansion meeting the above condition:
@@
Let $Q_{n}$ define a distribution over random functions $F_{in}(p)$, for any of the random functions listed in the Lemma, and define the function class $\mathcal{F}_{n}=\{F_{in}(\cdot)\to F_{in}(p):p\in\mathcal{S}\}$. Theorems 2.11.1 and Theorem 2.11.9 van der Vaart & Wellner (1996) imply weak convergence of the empirical process
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left(F_{in}(p)-f_{n}(p)-F_{in}(p^{\prime})-f_{n}(p^{\prime})\right)$
to a Gaussian process indexed by $p$ when for each $n$, the $\varepsilon$-bracketing number or $\varepsilon$-covering number of $\mathcal{F}_{n}$ under $Q_{n}$ is bounded by $(K/\varepsilon)^{V}$ for all $0<\varepsilon<K$, for finite constants $K,V>0$. Then, asymptotic equicontinuity follows from weak continuity, i.e. that for each $p\in\mathcal{S}$ and $p^{\prime}\in\mathcal{S}$, $\mathbb{E}[(F_{in}(p)-F_{in}(p^{\prime}))^{2}]$ is continuous in $p$. Weak continuity implies that the limiting Gaussian process has continuous sample paths.
@@
Now, pick any $c>0$. By asymptotic equicontinuity, $\limsup\limits_{n\rightarrow\infty}\mathbb{P}\left[\left\lvert R_{n}\right\rvert>c/\sqrt{n}\right]=0$. Furthermore
@@
where for the second inequality we used Cauchy-Schwarz. Since we have shown that $0\leq\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]\leq\epsilon$ for any $\epsilon>0$, we see that in fact $\limsup_{n\rightarrow\infty}n\,\mathbb{E}\left[R_{n}^{2}\right]=0$.
For all of the random functions listed, weak continuity holds by Assumption 5 and 6, and Lemma 21 provides the required subgaussian condition. A final check is htat
$\lim_{n\rightarrow\infty}\operatorname{Cov}\left[Z_{ij}(w,\,p+U_{i}),\ Z_{ij}(w,\,p^{\prime}+U_{i})\right]=\operatorname{Cov}\left[Z_{ij}(w,\,p),\ Z_{ij}(w,\,p^{\prime})\right],$ (91)
which is finite, and the same holds for $Y_{i}(w,p+U_{i})$ and $Y_{i}(W_{i},p+U_{i})$.